Line l goes through (-2,5) and is perpendicular to 6x-3y=1. Find the equation of line . Write the answer in standard form with integral coefficient with a positive coefficient for x.
Answer by richwmiller(17219)

-a/b=m
-6/-3=2
5=-1/2*2+b
5=-1+b
b=6
y=-x/2+6
x/2+y=6
x+2y=12